The point (-3,y) is on the line represented by the equation 26x-10y=100. Find the value of y for this point and then write the equation of the line in point-slope form using this point.
Answer by LinnW(1048)

26x-10y=100
add -26x to each side
-10y=-26x+100
multiply each side by -1
10y = 26x -100
divide each side by 10
y = (26/10)x -10
Using the pair (-3,y), set x = -3
y = (26/10)(-3) -10
y = -78/10 -100/10
y = -78/10 -100/10
y = -178/10
y = -17.8
substituting in the original equation,
26x-10y=100
26(-3)-10(-17.8)?=100
-78 + 178 = 100 This checks, so
the value for y in the pair (-3,y) is -17.8